# Sections of Lagrangian fibrations on holomorphic symplectic manifolds

### v4

@@ 1.2 C-symplectic structures and degenerate twistor deformations
Theorem 1.3:  Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration on a holomorphic symplectic manifold $(M,\Omega)$, and $S:\;X{\>\longrightarrow\>}M$ a smooth ($C^{\infty}$) section. Denote by $\eta\in\Lambda^{2}(X)$ the form $-S^{*}(\Omega)$. Then $\Omega+\pi^{*}\eta$ is C-symplectic, and, moreover, the image of $S$ is holomorphic in the complex structure $I_{\eta}$ induced by the C-symplectic form $\Omega+\pi^{*}\eta$.
@@ 1.5 Multiple fibers and primitivity
The discriminant of $\pi$ is the set of critical values of $\pi$. By a theorem of Hwang and Oguiso ([HO]), the discriminant is a divisor in $X$. We say that the fibers of $\pi$ have a reduced component in codimension 1 if the fiber $\pi^{-1}(s)$ is not multiple outside of a codimension 2 subset of $X$. It was conjectured that in codimension 1, the fibers of any Lagrangian fibration are actually reduced. In an earlier version of this paper, we argued that the assumption of having a reduced component for the fibers in codimension 1 is sufficient to have a deformation admitting a bimeromorphic section.
@@
Another place where the primitivity of the general fiber is used is Subsection 5.3. In the first step of the proof of 5.3, we invoke the same argument as used in Section 3 to produce the section over a curve. Here we apply it to extend this section to a piecewise smooth multisection of $\pi:\;M\to{\mathbb{C}}P^{n}$, in order to prove that the Dolbeault current associated with the section of $\pi$ in a neighbourhood of a line vanishes.
In [KV], Kamenova and Verbitsky have proven that $\pi^{*}([H])$ is primitive in $H^{*}(M)$, assuming that $\pi:\;M\to{\mathbb{C}}P^{n}$ has no multiple fibers in codimension 1. However, it does not imply that $\pi^{*}([H])$ is primitive in $H^{*}(M_{1})$. Indeed, in [Hel] I. Hellmann has shown that the Mukai system $M$ of rank two and genus two on a K3 surface admits a Lagrangian fibration $\pi:\;M\to X$ with a (very special) fiber of multiplicity 2. Therefore, $\pi$ has no bimeromorphic sections. If $\pi^{*}([H])$ were primitive in $H^{*}(M_{1})$, 1.1 would imply that $\pi$ has no multiple fibers.
@@
However, it is impossible to extend $\sigma:\;C\to X$ to a holomorphic section over the tubular neighbourhood of $C$. A smooth section, which necessarily exists, would violate 5.3, that is, the corresponding Dolbeault current is non-zero; in particular, it is impossible to extend the section $\sigma$ to a rational section of $\pi:\;X\to Y$, for any degenerate twistor deformation of $X$.
@@ 2.1 C-symplectic structures
Claim 2.2: The eigenvalues of $I$ are ±$\pm\sqrt{-1}\>$. Moreover, $I$ is diagonalizable over ${\mathbb{C}}$.
@@
Remark 2.4: One can reconstruct $I$ from the space $V^{1,0}\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$. Indeed, take $V^{0,1}=\overline{V^{1,0}}$, and let $I$ act on $V^{0,1}$ as $\sqrt{-1}\>\operatorname{Id}$, and on $V^{0,1}$ as $-\sqrt{-1}\>\operatorname{Id}$. Since the operator $I\in\operatorname{End}(V\otimes_{\mathbb{R}}{\mathbb{C}})$ commutes with complex conjugation, it is real, that is, it preserves $V\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$.
Definition 2.5: An almost complex structure on a real $2n$-manifold $M$ is an operator $I\in\operatorname{End}(TM)$ satisfying $I^{2}=-\operatorname{Id}_{TM}$, or equivalently, an $n$-dimensional subbundle $T^{1,0}M\subset TM\otimes_{\mathbb{R}}{\mathbb{C}}$ such that $T^{1,0}M\cap\overline{T^{1,0}M}=0$. The almost complex structure is called integrable (and $M$ is a complex manifold) if $T^{1,0}M$ satisfies $[T^{1,0}M,T^{1,0}M]\subset T^{1,0}M$.
@@
Proof: The rank of $\Omega$ is $2n$ because $\Omega^{n+1}=0$ and $\Omega^{n}\neq 0$ everywhere. Then $\ker\Omega\oplus\overline{\ker\Omega}=T_{\mathbb{C}}M$. The relation $[T^{0,1}_{\Omega}(M),T^{0,1}_{\Omega}(M)]\subset T^{0,1}_{\Omega}(M)$ follows from Cartan’s formula
@@ 2.3 Degenerate twistor deformations
2.4 Néron models
We introduce Néron models, starting from the traditional, scheme-theoretic definition (see e. g. [BLR, BHPT]).
The notion of Néron model can be defined in bigger generality, but for most practical purposes, people consider the Néron models of abelian fibrations. In this context, “abelian fibration” is understood as a proper holomorphic map $\pi:\;M{\>\longrightarrow\>}X$ equipped with a relatively ample bundle with an abelian variety as a general fiber and a section. If $X$ is quasiprojective, $\pi$ defines an abelian variety over the field of rational functions on $X$, and this section represents its zero.
Note that the the abelian fibration, as defined above, is equipped with a section, hence it has no multiple fibers.
Definition 2.23: Let $R$ be a Dedekind domain, $K$ its fraction field, and $X$ a scheme over $K$. For a scheme ${\cal Y}$ over $R$ we denote the corresponding scheme ${\cal Y}\times_{\operatorname{Spec}R}\operatorname{Spec}K$ over the fraction field by ${\cal Y}_{K}$. A Néron model of $X$ over $R$ is a smooth, separated finite-type $R$-scheme ${\cal X}$ such that ${\cal X}_{K}=X$ and ${\cal X}$ satisfies the Néron mapping property (NMP): for any smooth $R$-scheme ${\cal Y}$ and any $K$-morphism $f_{K}:\;{\cal Y}_{K}\to{\cal X}_{K}$, $f_{K}$ can be uniquely extended to a morphism $f:\;{\cal Y}{\>\longrightarrow\>}{\cal X}$.
Remark 2.24:  Uniqueness of a Néron model is clear. As follows from [BLR, Theorem 1.4/3], for any abelian variety over $K$, the Néron model always exists. Also, the Néron mapping property implies that the Néron model of an algebraic group over $K$ is a group scheme over $R$.
Remark 2.25: The Néron model was defined over a base which is a spectrum of a Dedekind domain. In the complex geometric setting, this amounts to defining a Néron model over the spectrum of a ring of germs of a smooth complex curve (which is a discrete valuation ring, that is, a local Dedekind domain). This is a local version of this notion, which can be globalized as follows. Define the Néron model of an abelian fibration $\pi:\;Y{\>\longrightarrow\>}C$ over a smooth complex curve $C$ as a complex variety ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ bimeromorphic to $Y$ which gives its Néron model if tensored by the ring ${\cal O}_{C,c}$ of germs of any point $c\in C$.
Note, however, that the Lagrangian fibrations we deal with do not, generally speaking, admit sections, hence they do not define abelian fibrations. If we restrict the fibration to a sufficiently small open set in the base, local sections exist, and we can always define its Néron model.
Theorem 2.26:  Let ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ be an abelian fibration over a smooth complex curve $C$. Assume that ${\cal Y}$ is smooth, algebraic, projective over $C$, and the relative canonical bundle of ${\cal Y}/C$ is trivial. Then the Néron model of ${\cal Y}_{K}$ is isomorphic to the smooth locus of $\varphi$, which is an open subset in ${\cal Y}$.
Proof: In an earlier version of this paper, we followed [BHPT], Section 3 and [HN, Proposition 5.1.8], referring to the notion of a Kulikov model. However, János Kollár [Ko2] pointed out that the definitions of the Kulikov model in these papers are different, and, for the proof, [BHPT] refers to a claim in [HN], which does not exist.
In the present version, we use a result of [Ko2]. Recall that a model of ${\cal Y}$ is a variety, fibered over $C$, bimeromorphic to ${\cal Y}$ over $C$, and with generic fiber isomorphic to that of ${\cal Y}$. A minimal model of an abelian fibration is a model which has terminal singularities and numerically trivial relative canonical bundle. Different flavors of this notion are called “Kulikov model” in [BHPT] and [HN, Proposition 5.1.8]. By [Ko2, Corollary 2 (2.2a)], the natural map from the smooth locus of a minimal model to the Néron model of ${\cal Y}$ is an isomorphism. 2.4 is a special case of this theorem.
This result immediately brings the following corollary. It is essentially the only result from this subsection which we are going to use.
Corollary 2.27:  Let ${\cal Y}\stackrel{{\scriptstyle\varphi}}{{{\>\longrightarrow\>}}}C$ be an abelian fibration over a smooth complex curve $C$ admitting local sections. Assume that ${\cal Y}$ is smooth, algebraic, and the relative canonical bundle of ${\cal Y}/C$ is pullback of a line bundle on $C$. Then the smooth locus of $\varphi$ is equipped with a structure of an algebraic group scheme over $C$.
Proof: Locally in $C$, the bundle ${\cal Y}/C$ is trivial, and therefore the smooth locus of $\varphi$ is the Néron model of ${\cal Y}$ (2.4). The Néron model is a group scheme by 2.4.
@@ 3.1 Constructing a smooth section over a curve using Néron models
Theorem 3.1:  Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$. Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology. Assume, moreover, that $\pi:\;\pi^{-1}(C)\to C$ has local sections. Then there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
Remark 3.2: The existence of local sections of $\pi:\;\pi^{-1}(C)\to C$ is needed because we use a Néron model of the fibration $\pi:\;\pi^{-1}(C)\to C$. In the next subsection we give another version of this theorem (3.2), which does not use Néron models and does not need this assumption.
Proof. Step 1: By Bertini’s theorem, the preimage $M_{1}:=\pi^{-1}(C)$ of $C$ is smooth. Denote by $Z\subset M_{1}$ the preimage of a general point. We are going to show that its fundamental class $[Z]\in H_{2}(M_{1})$ is primitive in the integral cohomology, that is, is not divisible by a non-invertible integer. Let $F:=\pi^{-1}(pt)$ be a general fiber of $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, and $[F]\in H_{2n}(M,{\mathbb{Z}})$ its homology class. Since $[F]$ is primitive, there exists a homology class $\alpha\in H_{2n}(M,{\mathbb{Z}})$ such that $[F]\cap\alpha=1$. Consider the class $\beta:=\alpha\cap[M_{1}]\in H_{2}(M_{1},{\mathbb{Z}})$; it can be realized as an intersection of $\alpha$ and $M_{1}$, if the chains defining $\alpha$, are transversal to $M_{1}$. By construction, this class intersects $[F]$ in one point, if counted with multiplicities, which implies that $[F]\cap\beta=1$.
@@
We are going to show that $z$ is independent of the choice of $e\in T$. Clearly, after identifying $T$ and $T_{0}$ by taking $e$ to be the origin, we obtain that $z=\sum_{i=1}^{k+1}x_{i}-\sum_{i=1}^{k}y_{i}$. However, (3.1) describes the dependence on the choice of $e$. Indeed, we can consider $T$ as a manifold equipped with a free and transitive action of $T_{0}$, which is independent of $e$; the expressions $(x_{i}-e)$ and $(y_{i}-e)$ belong to $T_{0}$, and $e+\sum_{i=1}^{k+1}(x_{i}-e)-\sum_{i=1}^{k}(y_{i}-e)$ denotes a point in $T$ shifted by these quantities. If we replace $e$ by $e+u$, the sum (3.1) would not change, hence it gives a section of the projection $\pi:\;M_{1}\to C$, defined outside of its discriminant locus.

### v5

@@ 1.2 C-symplectic structures and degenerate twistor deformations
Theorem 1.3: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration on a holomorphic symplectic manifold $(M,\Omega)$, and $S:\;X{\>\longrightarrow\>}M$ a smooth ($C^{\infty}$) section. Denote by $\eta\in\Lambda^{2}(X)$ the form $S^{*}(\Omega)$. Then $\Omega+\pi^{*}\eta$ is C-symplectic, and, moreover, the image of $S$ is holomorphic in the complex structure $I_{\eta}$ induced by the C-symplectic form $\Omega+\pi^{*}\eta$.
@@ 1.5 Multiple fibers and primitivity
The following result should be well-known, but we did not find it in the literature.
Lemma 1.7: Let $f:\;X{\>\longrightarrow\>}Y$ be a proper holomorphic map of complex manifolds, and $y\in Y$ any point. Then $f$ has a local section in a neighbourhood of $y$ if and only if the fiber $f^{-1}(y)$ has a component of multiplicity 1.
Proof: Assume that $f^{-1}(y)$ has a reduced component $Z$, and consider a transversal to $Z$ local submanifold $U$ of complimentary dimension intersecting $Z$ in a smooth point. Since $U\cap f^{-1}(y)=1$, for any $y^{\prime}\in Y$ sufficiently close to $y$, we also have $U\cap f^{-1}(y^{\prime})=1$. The implicit function theorem implies that the map $y^{\prime}\mapsto U\cap f^{-1}(y^{\prime})$ is a holomorphic map, defining a local section of $f$.
Conversely, assume that $f$ has a local section $s:\;V\to X$ in a neighbourhood $V\ni y$. Let $s(V)=U$; it is a smooth local submanifold of $X$. The intersection $U\cap f^{-1}(y)$ has multiplicity 1, hence it belongs to only one of the irreducible components $Z_{1}$ of $f^{-1}(y)$. However, unless $Z_{1}$ is smooth in the intersection point $U\cap Z_{1}$, the intersection $U\cap Z_{1}$ has multiplicity >1>1, which is impossible. Therefore, $Z_{1}$ is a smooth submanifold of $X$ in a neighbourhood of the point $U\cap Z_{1}$, and the multiplicity of this component is 1.
The discriminant of a proper holomorphic map $\pi$ is the set of critical values of $\pi$. By a theorem of Hwang and Oguiso ([HO]), the discriminant of a Lagrangian fibration is a divisor in $X$. We say that the fibers of $\pi$ have a reduced component in codimension 1 if the fiber $\pi^{-1}(s)$ is not multiple outside of a codimension 2 subset of $X$. It was conjectured that in codimension 1, the fibers of any Lagrangian fibration are actually reduced. In an earlier version of this paper, we argued that the assumption of having a reduced component for the fibers in codimension 1 is sufficient to have a deformation admitting a bimeromorphic section.
@@
Another place where the primitivity of the general fiber is used is Subsection 6.3. In the first step of the proof of 6.3, we invoke the same argument as used in Section 4 to produce the section over a curve. Here we apply it to extend this section to a piecewise smooth multisection of $\pi:\;M\to{\mathbb{C}}P^{n}$, in order to prove that the Dolbeault current associated with the section of $\pi$ in a neighbourhood of a line vanishes.
In [KV1], Kamenova and Verbitsky have proven that $\pi^{*}([H])$ is primitive in $H^{*}(M)$, assuming that $\pi:\;M\to{\mathbb{C}}P^{n}$ has no multiple fibers in codimension 1. However, it does not imply that $\pi^{*}([H])$ is primitive in $H^{*}(M_{1})$. Indeed, in [Hel] I. Hellmann has shown that the Mukai system $M$ of rank two and genus two on a K3 surface admits a Lagrangian fibration $\pi:\;M\to X$ with a (very special) fiber of multiplicity 2. Therefore, $\pi$ has no bimeromorphic sections. If $\pi^{*}([H])$ were primitive in $H^{*}(M_{1})$, 1.1 would imply that $\pi$ has no multiple fibers.
@@
However, it is impossible to extend $\sigma:\;C\to X$ to a holomorphic section over the tubular neighbourhood of $C$. A smooth section, which necessarily exists, would violate 6.3, that is, the corresponding Dolbeault current is non-zero; in particular, it is impossible to extend the section $\sigma$ to a rational section of $\pi:\;X\to Y$, for any degenerate twistor deformation of $X$.
@@ 2.1 C-symplectic structures
Claim 2.2: The eigenvalues of $I$ are $\pm\sqrt{-1}\>$. Moreover, $I$ is diagonalizable over $\mathbb{C}$}.
@@
Remark 2.4: One can reconstruct $I$ from the space $V^{1,0}\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$. Indeed, take $V^{0,1}=\overline{V^{1,0}}$, and let $I$ act on $V^{0,1}$ as $\sqrt{-1}\>\operatorname{Id}$, and on $V^{0,1}$ as $\sqrt{-1}\>\operatorname{Id}$. Since the operator $I\in\operatorname{End}(V\otimes_{\mathbb{R}}{\mathbb{C}})$ commutes with complex conjugation, it is real, that is, it preserves $V\subset V\otimes_{\mathbb{R}}{\mathbb{C}}$.
Definition 2.5: An almost complex structure on a real $2n$-manifold $M$ is an operator $I\in\operatorname{End}(TM)$ satisfying $I^{2}=-\operatorname{Id}_{TM}$, or equivalently, an $n$-dimensional subbundle $T^{1,0}M\subset TM\otimes_{\mathbb{R}}{\mathbb{C}}$ such that $T^{1,0}M\cap\overline{T^{1,0}M}=0$. The almost complex structure is called integrable (and $M$ is a complex manifold) if $T^{1,0}M$ satisfies $T^{1,0}M,T^{1,0}M]\subset T^{1,0}M$.
@@
Proof: The rank of $\Omega$ is $2n$ because $\Omega^{n+1}=0$ and $\Omega^{n}\neq 0$ everywhere. Then $\ker\Omega\oplus\overline{\ker\Omega}=T_{\mathbb{C}}M$. The relation $T^{0,1}_{\Omega}(M),T^{0,1}_{\Omega}(M)]\subset T^{0,1}_{\Omega}(M)$ follows from Cartan’s formula
@@ 2.3 Degenerate twistor deformations
3 Canonical models for Lagrangian fibrations
3.1 Smooth torsor bundle associated with a Lagrangian fibration
We relate some results about the Néron models and their generalizations to Lagrangian fibered hyperkähler manifolds. We follow [KV2]. The complex-geometric approach used in [KV2] is based on the notion of Kulikov model, discussed in [BLR, BHPT]; it was formally defined in [Ko2] as “canonical model”.
The main result which we use appeared as a part of Theorem 1.6 in [Ki]. In [KV2] we gave an independent proof of the following statement (see [Ki, Theorem 1.6]).
Theorem 3.1: Let $M$ be a holomorphic symplectic manifold, and $\pi:\;M\to X$ a Lagrangian fibration, that is, a proper holomorphic map which has holomorphic Lagrangian fibers and admits a relatively ample line bundle. Assume that $X$ is an open subset of a projective manifold, and $M_{i}\subset M$ the set of all points where the differental of $\pi$ has rank $\leqslant\dim X-i$. [^1] In [KV2], we show that $\operatorname{codim}\pi(M_{i})\geqslant i$. Denote by $X_{0}$ the set $X\backslash\pi(M_{2})$, and let $N\subset M$ be $\pi^{-1}(X_{0})\backslash M_{1}$. Then the natural projection $N\to X_{0}$ is equipped with a unique structure of a torsor of a smooth group scheme over the base $X_{0}$.
Proof: [KV2, Theorem 1.1].
Let us explain the terminology used in this statement.
Definition 3.2: Let $f:\;X_{0}\to Y$ be a smooth holomorphic submersion (not necessarily proper), and $s:\;Y\to X_{0}$ a section. Consider a map $X_{0}\times_{Y}X_{0}\stackrel{{\scriptstyle\mu}}{{\to}}X_{0}$ commuting with the projection to $Y$. We say that $(X_{0},s,\mu)$ is a holomorphic group bundle if the map $\mu$ is submersive and defines a group structure on each fiber $f^{-1}(y)$, for all $y\in Y$, and $s(y)$ is its identity element.
By definition, a torsor over an abelian group $G$ is a set where $G$ acts freely and transitively. However, the torsor structure can be defined without referring to an a priori given group structure as follows.
Definition 3.3: An abelian torsor is a set $S$ equipped with a map $\mu:\;S\times S\times S\to S$ such that
for any $e\in S$, the map $x\cdot y:=\mu(x,y,e)$ defines a commutative and associative product on $S$ with $e$ an identity element
for any fixed $y\in S$, the map $x\mapsto x\cdot y$ is bijective.
Remark 3.4: From the axioms of the abelian torsor, it is clear that $\cdot$ defines a structure of an abelian group on $S$.
Remark 3.5: The identity element axiom means that $\mu(x,y,y)=x$ for any $x$ and $y$. Given a group structure on $S$, the torsor operation $\mu$ can be defined as $\mu(x,y,z)=x+y-z$.
The most important property of abelian torsors is used many times throughout this paper.
Claim 3.6: Let $S$ be a set equipped with a structure of a torsor of an abelian group $G$, and $a_{1},...,a_{n},b_{1},...,b_{n+1}$ points in $S$. Consider $e\in S$, and let ++ denote the abelian group structure on $S$ associated with the unit $e$. Then the sum $\sum_{i=1}^{n+1}b_{i}-\sum_{i=1}^{n}a_{i}$ is independent from the choice of $e\in S$.
Proof: Let us denote the group operation on $S$ by $x+_{e}y:=\mu(x,e,y)$, and fix $u\in S$. Then $x+_{u}y=x+_{e}y-_{e}u$, and
$b_{1}+_{u}b_{2}+_{u}...+_{u}b_{n+1}=b_{1}+_{e}b_{2}+_{e}...+_{e}b_{n+1}-_{e}(n+1)u.$
Then
$b_{1}+_{u}b_{2}+_{u}...+_{u}b_{n+1}-_{u}a_{1}-_{u}a_{2}-_{u}...-_{u}a_{n}=$
$b_{1}+_{e}b_{2}+_{e}...+_{e}b_{n+1}-_{e}a_{1}-_{e}a_{2}-_{e}...-_{e}a_{n}-_{e}(n+1)u+_{e}nu=$
$b_{1}+_{e}b_{2}+_{e}...+_{e}b_{n+1}-_{e}a_{1}-_{e}a_{2}-_{e}...-_{e}a_{n}$
(see also the proof of 4.1, Step 2).
3.2 Néron models for Lagrangian fibrations
When a Lagrangian fibration $\pi:\;M\to X$ is equipped with a holomorphic section, the torsor bundle defined in 3.1 is actually a group bundle. In [Ki], this group bundle was interpreted as a part of the multi-dimensional Néron model of $M$. Throughout this paper, we often refer to this construction as to the Néron model construction for the Lagrangian fibration. Here is the relevant statement from [KV2] formulated for Lagrangian fibrations equipped with holomorphic sections.
Definition 3.7: Let $M_{i}$ be the set of all points $m\in M$ such that the differential $D\pi$ has rank $\leqslant\dim X-i$. The set $\pi(M_{1})$ of singular values of a Lagrangian fibration map $\pi:\;M\to X$ is called is discriminant, denoted by $\operatorname{Discr}\subset X$. Let $\operatorname{Discr}_{i}:=\pi(M_{i+1})$. Clearly, $\operatorname{Discr}=\operatorname{Discr}_{0}\supset\operatorname{Discr}_{1}\supset$.... By [KV2, Claim 2.9], $\operatorname{codim}_{\operatorname{Discr}}\operatorname{Discr}_{i}\geqslant i$.
Remark 3.8: Note that $\operatorname{Discr}_{i}$ does not necesarily coincide with the $i$-th singular stratum of $\operatorname{Discr}$. However, it has the same dimension.
Theorem 3.9: Let $M$ be a hyperkähler manifold, $f:\;M\to X$ a Lagrangian fibration, and $M_{0}\subset M$ its smooth locus. Denote by $J(f)$ the manifold $M_{0}\cap f^{-1}(X\backslash\operatorname{Discr}_{1})$. Assume that $f$ admits a holomorphic section $s:\;X\to M_{0}$. Then there exists a unique group bundle structure on $f:\;J(f)\to X\backslash\operatorname{Discr}_{1}$ with the following properties:
For each $y\in X\backslash\operatorname{Discr}_{1}$, $s(y)$ is the identity element of the fiber $f^{-1}\cap J(f)$.
[…]
Proof: [KV2, Theorem 2.13].
@@ 4.1 Constructing a smooth section over a curve using Néron models
Theorem 4.1: Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$. Assume that the general fiber $F$ of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in the integral homology. Assume, moreover, that the fibers of $\pi$ are reduced away from a codimension $\geqslant 2$ subset of $X={\mathbb{C}}P^{n}$. Then
there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
Moreover, this section can be chosen homologous to the homology class $\alpha\cap\pi^{-1}(C)\in H_{2}(\pi^{-1}(C),{\mathbb{Z}})$, where $\alpha\in H_{2n}(M,{\mathbb{Z}})$ a homology class such that $\alpha\cap[F]=1$, and $F]\in H_{2n}(M,{\mathbb{Z}})$ the homology class of a general fiber of $\pi$.
Remark 4.2: The existence of local sections of $\pi:\;\pi^{-1}(C)\to C$ is needed in 4.3 (ii); it is also used in the original proof of 4.3 (i) based on the Néron models of the fibration $\pi:\;\pi^{-1}(C)\to C$ (given below). In Subsection 4.3 we give another version of 4.3 (i), which does not use Néron models and does not need this assumption.
We end this subsection with a proof of 4.1 (i). In Subsection 4.2, we explain why 4.1 (ii) is non-trivial and prove it, and in Subsection 4.3, we give another proof of 4.1 (i).
Proof of 4.1 (i), step 1: By Bertini’s theorem, the preimage $M_{1}:=\pi^{-1}(C)$ of $C$ is smooth. Denote by $Z\subset M_{1}$ the preimage of a general point. We are going to show that its fundamental class $Z]\in H_{2}(M_{1})$ is primitive in the integral cohomology, that is, is not divisible by a non-invertible integer. Let $F:=\pi^{-1}(pt)$ be a general fiber of $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, and $F]\in H_{2n}(M,{\mathbb{Z}})$ its homology class. Since $[F]$ is primitive, there exists a homology class $\alpha\in H_{2n}(M,{\mathbb{Z}})$ such that $F]\cap\alpha=1$. Consider the class $\beta:=\alpha\cap[M_{1}]\in H_{2}(M_{1},{\mathbb{Z}})$; it can be realized as an intersection of $\alpha$ and $M_{1}$, if the chains defining $\alpha$, are transversal to $M_{1}$. By construction, this class intersects $[F]$ in one point, if counted with multiplicities, which implies that $F]\cap\beta=1$.
@@
We are going to show that $z$ is independent of the choice of $e\in T$. Clearly, after identifying $T$ and $T_{0}$ by taking $e$ to be the origin, we obtain that $z=\sum_{i=1}^{k+1}x_{i}-\sum_{i=1}^{k}y_{i}$. However, (4.1) describes the dependence on the choice of $e$. Indeed, we can consider $T$ as a manifold equipped with a free and transitive action of $T_{0}$, which is independent of $e$; the expressions $(x_{i}-e)$ and $(y_{i}-e)$ belong to $T_{0}$, and $e+\sum_{i=1}^{k+1}(x_{i}-e)-\sum_{i=1}^{k}(y_{i}-e)$ denotes a point in $T$ shifted by these quantities. If we replace $e$ by $e+u$, the sum (4.1) would not change, hence it gives a section of the projection $\pi:\;M_{1}\to C$, defined outside of its discriminant locus. For an abstract version of this argument (formulated for any abelian torsor) see 3.1.